OpenStudy (anonymous):

an automobile traveling 30.0 m/s has wheels of 0.750 m diameter. (a) what is the angular speed of the wheels about the axle? (b) if the car is brought to a stop uniformly in 10.0 s, what is the angular acceleration? (c) How many times do the wheels turn during thus braking period?

OpenStudy (anonymous):

\[a = -3m/s^2\]

OpenStudy (anonymous):

\[aT = \alpha r\]

OpenStudy (anonymous):

\[= -3 * 0.75/ 2 = -1.125\]

OpenStudy (anonymous):

\[s = ut + 1/2 a t^2\]

OpenStudy (anonymous):

\[= 300 - 1/2 * 1.25 * 10 * 10 = 243.75\]

OpenStudy (anonymous):

@shubhamsrg @ajprincess

OpenStudy (anonymous):

\[2 * 3.14 * 0.375 = 2.3\]

OpenStudy (anonymous):

= 2.355

OpenStudy (anonymous):

@sarahvince Do u have the answer key???

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

so the answer in letter c is 2,355?

OpenStudy (anonymous):

i am not sure of my answers

OpenStudy (anonymous):

No : of rotations = n * Circumfrence = Distance covered

OpenStudy (anonymous):

im confused hehehe

OpenStudy (anonymous):

Do u knw abt Circular Motion??

OpenStudy (anonymous):

yes thats our latest lesson

OpenStudy (anonymous):

Do u knw angular velocity , centripetal acceleartion

OpenStudy (anonymous):

i just want to makes ure my answers are correct. hehehe

OpenStudy (anonymous):

wat is ur answers

OpenStudy (anonymous):

@shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg @shubhamsrg

OpenStudy (anonymous):

@dpaInc @lgbasallote @UnkleRhaukus @hba

OpenStudy (anonymous):

@maheshmeghwal9

OpenStudy (anonymous):

@ujjwal @ujjwal @ujjwal @ujjwal @ujjwal @ujjwal

OpenStudy (anonymous):

@kartiksriramk @kartiksriramk @kartiksriramk @kartiksriramk @kartiksriramk @kartiksriramk @kartiksriramk

OpenStudy (anonymous):

@Compassionate @experimentX

OpenStudy (anonymous):

i think there is wrong in my answer can u correct it @experimentX

OpenStudy (experimentx):

v = wr

OpenStudy (anonymous):

that is linear speed

OpenStudy (experimentx):